OpenStudy (anonymous):

"write an equation for the linear function f satisfying the given conditions" f(-3)=8 and f(1)=4

OpenStudy (anonymous):

f(-3)=12

OpenStudy (anonymous):

I need to know how to show my work for this type of problem

OpenStudy (anonymous):

im in algebra 2 part 1 so.... do you want to put it into y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok take one problem f(-3)=8 im going to solve for x first

OpenStudy (anonymous):

\[f(-3)=8\] add 3 to both sides x=11

OpenStudy (anonymous):

solve for y f(1)=4 subtract 1 from both sides y=3

OpenStudy (anonymous):

plug x and y in to one equation

OpenStudy (anonymous):

f(y)=mx+b f(3)=1(11)+4 f(3)=7 subtract 3 from both sides slope is 4

OpenStudy (anonymous):

then its simple y=mx+b 3=4(11)+4

OpenStudy (anonymous):

3=48?

OpenStudy (anonymous):

im confused. I have to write an equation using the above f(-3)=8 and f(1)=4

OpenStudy (anonymous):

Can you find two points using f(-3)=8 and f(1)=4?

OpenStudy (anonymous):

(-3 , 1 ) ?

OpenStudy (anonymous):

For f(-3)=8, there is an x of -3 and a y of 8 --> (-3,8) For f(1)=4, there is an x of 1 and a y of 4 --> (1,4)

OpenStudy (anonymous):

Oh ok, so now that I know my x's and y's how would I set that up as an equation showing my work

OpenStudy (anonymous):

Use the equation \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] to find your slope and then you could either use y=mx+b or point-slope formula

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.